OpenStudy (anonymous):

say i have 4/7x =40/77 and i have 2 clear them how do i do this

OpenStudy (id21):

\[4\times77=7x \times 40;1\times11=x \times 10;x=11/10=1,1\]

OpenStudy (anonymous):

so the eqation is really 1

OpenStudy (id21):

You can divide both sides by 4 and multiply both sides by 7 and you will get \[(4\times7)/(7x \times4)=(40\times 7)/(77 \times 4) -> 1/x=10/11\]